# Analyzing Key Objectives in Human-to-Robot Retargeting for Dexterous Manipulation

### v1

@@ Abstract
Index Terms:
Dexterous manipulation, multi-fingered hand, human-to-robot retargeting, teleoperation.
@@ I Introduction
Transferring dexterity from humans to robots is a promising field in dexterous manipulation research, as the complexity of dexterous manipulation tasks poses challenges for classical analytical approaches [1, 2, 3, 4]. In either collecting robot demonstrations through human teleoperation or learning from offline human manipulations, an essential component is kinematic retargeting, which refers to kinematically translating the human hand configuration to robot hand joint positions. One key challenge of retargeting is that, due to the differences in morphology and degrees of freedom (DoFs) between human hands and current robot hands, the human hand configurations can not be exactly reproduced by the robot hand. Consequently, it is inevitable to focus on only partial configuration of the human hand that are more crucial to manipulation tasks when designing the retargeting algorithms.
@@
The most straightforward retargeting approach is direct joint-to-joint mapping, where the robot hand joints follow a manually defined transformation of the human hand joints [5, 6, 7, 8, 9, 10, 11, 12]. These approaches require considerable manual efforts to define the joint-space transformation for a certain robot hand. Task-space retargeting is used more widely, which relies on inverse kinematics (IK) solving to determine robot hand joint configuration based on human hand keypoints. The global fingertip position is the fundamental retargeting objective to align human and robot hand in task-space [13, 14, 15, 16, 17]. In most works, the relative fingertip position to the wrist serves as the retargeting objective [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Distance between two set of vectors from the wrist to the fingertips is minimized to obtain the robot’s optimal joint configurations, thereby preserving the overall hand shape. Some works include finger orientations in the objective to preserve more human-like shapes. The finger shapes are represented in the form of finger bending information by including a set of vectors from palm to middle phalanx [19]. More explicitly, the directions of five proximal phalanges and distal phalanges [30] are used to represent orientations of the finger-roots and fingertips [31]. Additionally, relative position among fingertips can be included to address the challenge of pinch grasps [28], in which a switching weight is used to prioritize the fingers involved in pinching, and the corresponding reference distance on human hand is adjusted to a minimal distance to encourage fingertip contact.
These retargeting approaches have been deployed in real-world manipulation, but they differ in their choices of objectives. Due to the lack of comparative studies, it remains unclear which objectives are more dispensable for which types of manipulation tasks and whether conflicts exist between them. A previous survey [32] reviews and classifies the common retargeting methods, but it does not focus on retargeting objectives and not provide experimental evaluations.
@@ II Method
$\mathcal{L}_{\text{fingertip\_pos}}=\sum_{i=1}^{N}\left\|\bm{v}_{i}^{h}-\bm{v}_{i}^{r}\right\|_{2}^{2},$ (2)
where $\bm{v}_{i}$ is the vector from the wrist to the $i^{th}$ fingertip on human and robot hand, and $N$ is the number of fingers. Note that this term needs to be coordinated with the following pinch objective, for which the details are provided in Appendix.
Relative position among fingertips: We use vectors from the thumb to primary fingers (index, middle, ring) to represent the relative positions among fingertips, which is crucial for pinching. We adopt a similar formulation to DexPilot [28] with a switching weight function $s(d_{i})$ and a distance rescaling function $l(d_{i})$:
@@
where $\bm{\gamma}_{i}$ is the vector from the thumb fingertip to the fingertip of the $i^{th}$ primary finger, $d_{i}$ is the length of $\bm{\gamma}_{i}$ and $\hat{\bm{\gamma}}_{i}=\frac{\bm{\gamma}_{i}}{d_{i}}$. Instead of using a discrete weight function as [28], we use continuous weight and rescaling functions to ensure smooth transitions. Specifically, our rescaling function linearly rescales human fingertip distances from $[\epsilon_{2},\epsilon_{1}]$ to $[0,\epsilon_{1}]$. The details are provided in Appendix.
Fingertip orientations: To represent fingertip orientations, we include another set of vectors. In contrast with methods such as DexMV [19] that define the vectors from the wrist to the middle phalanx, our formulation defines the vectors from the distal interphalangeal (DIP) joints to the fingertips, which represent fingertip orientations more directly:
@@
where two joint-space regularization terms are additionally considered. The joint position regularization $\mathcal{L}_{\text{joint}}=\sum_{j=1}^{m}w_{j}^{\rm pos}\left\|q_{j}-\bar{q}_{j}\right\|_{2}^{2}$ penalizes the deviation of some joints from a pre-defined joint configuration $\bar{\bm{q}}$. The joint velocity regularization $\mathcal{L}_{\text{vel}}=\sum_{j=1}^{m}w_{j}^{\rm vel}\left\|q_{j}-q_{j}^{\text{prev}}\right\|_{2}^{2}$ penalizes large changes in joint positions compared to previous timestep to encourage trajectory smoothness.
@@ III-A Evaluation Setup
Implementation: Simulation studies involve a Leap Hand [33] and a shadow hand, both mounted on a Franka Emika Panda arm. Real-world experiments are conducted on a teleoperation system with a Leap Hand, a Panda arm, and an Apple Vision Pro for human hand detection [34]. The retargeting objective is optimized in real time using the Sequential Least-Squares Quadratic Programming in the NLopt library [35].
@@
  1. 1.
    Pick up a box, rotate it by 90 degrees, and place it down, which represents commonly used pick-and-place tasks in dexterous teleoperation.
  2. 2.
@@
  3. 3.
    Pick up five upright-standing screws of different sizes and place them on target positions vertically, where precise pinches are important.
For each task, we sequentially evaluate the full retargeting objective and the seven ablations (A1 to A7), and repeat the entire sequence three times. The above tests are conducted twice with two human pilots. For task 1 and 2, we use mean completion time to assess performance of pilots, as all trials are successful. For task 3, we use success rate for comparison. The manipulation results are summarized in Fig. 4.
@@ III-B Analysis of Results
IV Conclusion
This work analyzes the significance of different objectives in human-to-robot retargeting for dexterous manipulation through both kinematic posture retargeting and three representative real-world manipulation tasks. The comprehensive results demonstrate that 1) the fingertip pinch objective is crucial for manipulation tasks involving precise fingertip coordination; 2) the fingertip orientation objective should be included for tasks sensitive to finger orientation rather than solely position; 3) allowing wrist pose adjustment instead of using exact human wrist pose benefits the accuracy of fingertip poses; 4) joint position regularization makes the retargeting postures appear more natural while having negligible negative impacts; and 5) all terms using the proposed formulation do not demonstrate conflicts and perform well in all tasks. We believe this study provides valuable insight for designing retargeting algorithms in future work on learning dexterous manipulation from humans or dexterous teleoperation.
@@ IV Conclusion
Appendix Supplementary results and implementation details are provided in Appendix, which includes hyper-parameters of the retargeting algorithm, the detailed formulation of the pinch objective, and additional quantitative results of kinematic posture retargeting on another trajectory involving finger crossing motions and on the Shadow Hand. The appendix is also available on the project website.

### v2

@@ I Introduction
Transferring dexterity from humans to robots is a promising field in dexterous manipulation research, as the complexity of dexterous manipulation tasks poses challenges for classical analytical approaches [pang2023global, jiang2024contact, yu2025robotic, yu2024hand]. In either collecting robot demonstrations through human teleoperation or learning from offline human manipulations, an essential component is kinematic retargeting, which refers to kinematically translating the human hand configuration to robot hand joint positions. One key challenge of retargeting is that, due to the differences in morphology and degrees of freedom (DoFs) between human hands and current robot hands, the human hand configurations can not be exactly reproduced by the robot hand. Consequently, it is inevitable to focus on only partial configuration of the human hand that are more crucial to manipulation tasks when designing the retargeting algorithms.
@@
The most straightforward retargeting approach is direct joint-to-joint mapping, where the robot hand joints follow a manually defined transformation of the human hand joints [gao2023hand, huang2025dih, wei2023adaptive, guo2024telephantom, liarokapis2013telemanipulation, giudici2023feeling, arunachalam2023holo, iyer2024open]. These approaches require considerable manual efforts to define the joint-space transformation for a certain robot hand. Task-space retargeting is used more widely, which relies on inverse kinematics (IK) solving to determine robot hand joint configuration based on human hand keypoints. The global fingertip position is the fundamental retargeting objective to align human and robot hand in task-space [wang2024dexcap, shaw2024bimanual, zhang2025doglove, arunachalam2023dexterous, chen2024arcap]. Most works use vectors from wrist to fingertips as the retargeting objective and minimize their discrepancy between human and robot hands to preserve overall hand shape[antotsiou2018task, qin2022dexmv, qin2023anyteleop, ding2024bunny, yang2024ace, huang2024designing, shaw2024learning, yuan2024cross, cheng2024open, romero2024eyesight, handa2020dexpilot, sivakumar2022robotic]. Some works further incorporate fingertip orientation into the objective, using vectors from the palm to the middle phalanx to represent finger bending information[qin2022dexmv], or using the directions of proximal phalanges and distal phalanges [enwiki:1275529475] to represent orientations of the finger-roots and fingertips [li2019vision]. Additionally, relative position among fingertips can be included to address the challenge of pinch grasps [handa2020dexpilot], in which a switching weight is used to prioritize the fingers involved in pinching, and the corresponding reference distance on human hand is adjusted to a minimal distance to encourage fingertip contact.
These retargeting approaches have been deployed in real-world manipulation, but they differ in their choices of objectives. Due to the lack of comparative studies, it remains unclear which objectives are more dispensable for which types of manipulation tasks and whether conflicts exist between them. A previous survey [meattini2022human] reviews and classifies the common retargeting methods, but it does not focus on retargeting objectives and not provide experimental evaluations.
@@ II Method
$\mathcal{L}_{\text{fingertip\_pos}}=\sum_{i=1}^{N}\tilde{s}(d_{i})\left\|\bm{v}_{i}^{r}-\bm{v}_{i}^{h}\right\|^{2},$ (2)
where $\bm{v}_{i}$ is the vector from the wrist to the $i^{th}$ fingertip on human and robot hand, and $N$ is the number of fingers. $\tilde{s}(d_{i})$ is a switching weight function to balance fingertip positions relative to the wrist and the thumb, where
$\tilde{s}(d_{i})=\operatorname{sigmoid}(d_{i},\epsilon_{1},-10),$
This term is designed to be coordinated with the following pinch objective $\mathcal{L}_{\text{pinch}}$, so that the sum of pinch term switching weight $s(d_{i})$ and the switching weight here $\tilde{s}(d_{i})$ will be a fixed number. Later, in ablation studies where the pinch term is removed (A1, A6 and A8), we set $\tilde{s}(d_{i})$ to be a constant $1.0$.
Relative position among fingertips: We use vectors from the thumb to primary fingers (index, middle, ring) to represent the relative positions among fingertips, which is crucial for pinching. We adopt a similar formulation to DexPilot [handa2020dexpilot] with a switching weight function $s(d_{i})$ and a distance rescaling function $l(d_{i})$:
@@
where $\bm{\gamma}_{i}$ is the vector from the thumb fingertip to the fingertip of the $i^{th}$ primary finger, $d_{i}=\left\|\bm{\gamma}_{i}^{h}\right\|$ and $\hat{\bm{\gamma}}_{i}^{h}=\frac{\bm{\gamma}_{i}^{h}}{d_{i}}$. Instead of using a discrete weight function as DexPilot, we use a continuous weight function to ensure smooth transitions:
$s(d_{i})=\operatorname{sigmoid}(d_{i},\epsilon_{1},10),$
where $\operatorname{sigmoid}(\cdot)$ is the sigmoid function defined as follows:
$\operatorname{sigmoid(x,c,w)}=\frac{1}{1+e^{w(x-c)}}.$
Our distance rescaling function is defined as follows:
$l(d_{i})=\begin{cases}0,&d_{i}<\epsilon_{2}\\ \frac{\epsilon_{1}}{\epsilon_{1}-\epsilon_{2}}(d_{i}-\epsilon_{2}),&\epsilon_{2}\leq d_{i}\leq\epsilon_{1}\\ d_{i},&d_{i}>\epsilon_{1},\end{cases}$ (4)
where fingertip distance within pinching range $\epsilon_{2},\epsilon_{1}]$ is linearly rescaled into $0,\epsilon_{1}]$. This ensures a continuous transition in the pinching range and avoids sudden changes around the threshold $\epsilon_{1}$. In practice we set $\epsilon_{1}=1\times 10^{-1}\,\text{m}$ and $\epsilon_{2}=1\times 10^{-2}\,\text{m}$.
Fingertip orientations: To represent fingertip orientations, we include another set of vectors. In contrast with methods such as DexMV [qin2022dexmv] that define the vectors from the wrist to the middle phalanx, our formulation defines the vectors from the distal interphalangeal (DIP) joints to the fingertips, which represent fingertip orientations more directly:
@@
where two joint-space regularization terms are additionally considered. The joint position regularization $\mathcal{L}_{\text{joint}}=\sum_{j=1}^{m}w_{j}^{\rm pos}\left\|q_{j}-\bar{q}_{j}\right\|_{2}^{2}$ penalizes the deviation of some joints from a pre-defined joint configuration $\bar{\bm{q}}$. The joint velocity regularization $\mathcal{L}_{\text{vel}}=\sum_{j=1}^{m}w_{j}^{\rm vel}\left\|q_{j}-q_{j}^{\text{prev}}\right\|_{2}^{2}$ penalizes large changes in joint positions compared to previous timestep to encourage trajectory smoothness. Its effect is illustrated by the joint position/velocity/acceleration profiles in Appendix.
@@ III-A Evaluation Setup
Implementation: Simulation studies involve a Leap Hand [shaw2023leaphand] and a Shadow Hand, both mounted on a Franka Emika Panda arm. Real-world experiments are conducted on a teleoperation system with a Leap Hand, a Panda arm, and an Apple Vision Pro for human hand detection [park2024avp]. The retargeting objective is optimized in real time using the Sequential Least-Squares Quadratic Programming in the NLopt library [NLopt].
In the real-world experiments, we use joint-space position control for both the Panda arm and the robot hand. At each retargeting step (20 Hz), the retargeting optimizer outputs a desired joint configuration $\bm{q}_{t}$ for all actuated joints. This configuration is first upsampled to a 100 Hz command stream via interpolation and then sent to the hardware. The LEAP Hand executes the control command via a built-in current-based PD controller. The Franka arm runs an impedance controller at a higher 1000 Hz with strict acceleration and jerk limits, so the 100 Hz joint targets are further interpolated to match the 1000 Hz control frequency before being executed by the arm controller.
In Appendix, we further discuss the runtime and latency of our system, as well as its robustness to tracking noise and occlusions.
@@
  1. 1)
    Pick up a box, rotate it, and place it down, representing common pick-and-place tasks in dexterous teleoperation.
  2. 2)
@@
  3. 3)
    Pick up five different upright-standing screws and place them vertically, where precise pinches are important.
For each task, we sequentially evaluate the full retargeting objective and the seven ablations (A1 to A7), and repeat the entire sequence three times. For task 1 and 2, we use mean completion time to assess performance of pilots, as all trials are successful. For task 3, we use success rate for comparison. The manipulation results are summarized in Fig. 4.
@@ III-B Analysis of Results
III-C Comparison with Existing Retargeting Approaches
Figure 5: (a) Results of kinematic posture retargeting on finger pinch trajectories. (b) Results of the real-world manipulations. Task 1 and 2 are assessed by task time, while Task 3 is evaluated by success rate. The bars represent the full objective, two ablations (A2, A4), DexMV, and DexPilot.
In addition to the ablation studies, we also compare our full proposed method with two representative retargeting approaches, DexMV and DexPilot [qin2022dexmv, handa2020dexpilot]. In our ablation setting, A2 and A4 are designed to closely mimic the key design choices of DexPilot and DexMV respectively. Specifically, A2 uses the raw fingertip pinch distance in the pinch term; unlike DexPilot, it does not include the fingertip distance rescaling function to ensure minimal distance between the thumb and a primary finger and force minimum separation distance between two primary fingers. A4 replaces the vectors from DIP joints to fingertips with vectors from the wrist to DIP joints, which is similar to the fingertip orientation formulation in DexMV (we use DIP joints instead of the PIP joints used in the original DexMV).
In contrast, the pinch term in our full objective differs from DexPilot in two aspects. First, we use a continuous switching weight $s(d_{i})$ based on a sigmoid function, instead of the discrete weights in DexPilot. Second, our distance rescaling function $l(d_{i})$ varies continuously on $\epsilon_{2},\epsilon_{1}]$, while DexPilot directly clamps fingertip distances below a fixed threshold. As a result, in our objective, both the rescaled distance and its weight change smoothly as the human pinch distance crosses threshold $\epsilon_{2}$ and $\epsilon_{1}$, which leads to smoother and more stable motions during precise pinching in real-world teleoperation.
To further ensure a fair comparison beyond these approximate variants, we additionally implement DexMV and DexPilot by directly following their objective functions and hyper-parameter settings. For DexPilot, we follow the code in the dex-retargeting repository [qin2023anyteleop]. Since their formulations are defined only for hand retargeting, we integrate them into our arm-hand retargeting pipeline by first applying their hand retargeting objective in the wrist frame and then separately solving the arm IK to track the resulting wrist pose.
We evaluate the full objective, the two ablations (A2 and A4), DexMV, and DexPilot on the LEAP-hand kinematic retargeting trajectories on finger pinch motions described in Sec. III-A, and compute the same four metrics as in Fig. 3: fingertip global position error, fingertip relative position to the wrist, fingertip relative position to the thumb, and fingertip orientation error. We also evaluate them on the same three real-world manipulation teleoperation tasks as Sec. III-A.
Figure 5 summarizes the quantitative kinematic posture retargeting and real-world manipulation results. The results show that, on the kinematic retargeting trajectories (Fig. 5(a)), the full objective consistently achieves lower fingertip global position error, fingertip relative position to the thumb, and fingertip orientation error compared to DexMV and DexPilot.
The real-world teleoperation experiments in Fig. 5(b) further validate our design choice. Across all three manipulation tasks, the full objective achieves the shortest or comparable task times and the highest success rates for both human pilots. DexMV and DexPilot, by contrast, generally lead to slower executions and more failures, especially on Task 3 which requires precise fingertip pinch. In particular, DexMV fails in every trial of Task 3 because it does not include the fingertip pinch term and therefore cannot close the gap between thumb and index fingertips, which is similar to the results of ablation A1 where we remove the pinch term. These results demonstrate that the proposed objective not only improves kinematic retargeting quality but also leads to more efficient and reliable performance in real-world dexterous teleoperation compared to existing retargeting approaches.
Supplementary materials are listed in Table II.
@@ III-C Comparison with Existing Retargeting Approaches
Appendix Supplementary results and implementation details are provided in Appendix, which includes additional quantitative results of kinematic posture retargeting on another trajectory involving finger crossing motions and on the Shadow Hand, explanation of the joint velocity regularization term, and implementation details such as hyper-parameter settings and discussion on runtime, latency, and robustness. The Appendix is also available on the project website.
@@
IV Conclusion
This work analyzes the significance of different objectives in human-to-robot retargeting for dexterous manipulation through both kinematic posture retargeting and three representative real-world manipulation tasks. The comprehensive results demonstrate that 1) the fingertip pinch objective is crucial for manipulation tasks involving precise fingertip coordination; 2) the fingertip orientation objective should be included for tasks sensitive to finger orientation rather than solely position; 3) allowing wrist pose adjustment instead of using exact human wrist pose benefits the accuracy of fingertip poses; 4) joint position regularization makes the retargeting postures appear more natural while having negligible negative impacts; and 5) all terms using the proposed formulation do not demonstrate conflicts and perform well in all tasks. We also show in the experiments that the complete objective we purpose leads to more accurate and reliable performance than existing retargeting approaches. We believe this study provides valuable insight for designing retargeting algorithms in future work on learning dexterous manipulation from humans or dexterous teleoperation. In future work, we plan to extend the real-world evaluation to more teleoperators and tasks to better capture user variability and validate the generality of the proposed retargeting formulation.
Figure 6: Additional results on kinematic posture Retargeting. (a) Kinematic posture retargeting results using Leap hand on another trajectory involving finger crossing motion. (b) Kinematic posture retargeting results using Shadow hand on the same pinch motion trajectory as in the main text. (c) Kinematic posture retargeting results using Shadow hand on the trajectory involving finger crossing motion.
Additional Kinematic Posture Retargeting Results
Three sets of additional quantitative results of kinematic posture retargeting on another trajectory involving finger crossing motions and on the Shadow Hand are shown in Fig. 6. Similar conclusions to the main text can be derived.
Explanation of Joint Velocity Regularization Term
As introduced in Sec. II, the joint-velocity regularization term
$\mathcal{L}_{\text{vel}}=\sum_{j=1}^{m}w_{j}^{\rm vel}\left\|q_{j}-q_{j}^{\text{prev}}\right\|_{2}^{2}$
penalizes large changes in joint positions between consecutive timesteps and is intended to encourage smooth retargeted joint trajectories.
To illustrate the effect of this term, we compare the full objective in (6) with an ablation that removes $\mathcal{L}_{\text{vel}}$. We use a real-world teleoperation trajectory that involves finger-crossing motions with the index finger and generate the corresponding retargeted joint trajectories under both settings. The trajectories are evaluated on the three joints of the index finger (MCP, PIP, and DIP) on the LEAP Hand.
Figure 7: Index MCP/PIP/DIP joint position/velocity/acceleration profiles with and without $\mathcal{L}_{\text{vel}}$. For all three joints, the position trajectories are smooth and free of oscillations in both cases. However, removing $\mathcal{L}_{\text{vel}}$ consistently leads to larger peaks in the velocity and especially in the acceleration profiles, indicating sharper changes in joint motion during the finger-crossing motion.
Figure 7 shows the joint position, velocity, and acceleration profiles for the index MCP, PIP, and DIP joints, respectively, under the full objective (blue line) and the variant without $\mathcal{L}_{\text{vel}}$ (orange line). These results empirically support the importance of $\mathcal{L}_{\text{vel}}$ in suppressing high-frequency variations and reducing acceleration spikes, thereby improving the overall trajectory smoothness.
Implementation
Hyper-parameter Setting
The hyper-parameters used in the retargeting objective are listed in Table III.
TABLE III: Hyper-parameters
Hyper Parameter Value
$\lambda_{1}$ 10
$\lambda_{2}$ 0.1
$\lambda_{3}$ 1
$\lambda_{4}$ 10
$\lambda_{5}$ 10
$w_{j}^{\text{pos}}$ (Leap hand) 0.5, $j=7,11,15,18$
0.1, $j=20$
0, else
$w_{j}^{\text{pos}}$ (Shadow Hand) 0.5, $j=9,13,17,22$
0.1, $j=26$
0, else
$w_{j}^{\text{vel}}$ (Leap hand) 0.1, $j=0\sim 6$
0.01, $j=7\sim 22$
$w_{j}^{\text{vel}}$ (Shadow Hand) 0.1, $j=0\sim 6$
0.01, $j=7\sim 30$
The weights $\lambda_{1}$$\lambda_{5}$ reflect the relative importance of different retargeting objectives. We assign larger weights of 10 to the thumb-tip position, fingertip orientation, and pinch terms ($\lambda_{1}$, $\lambda_{4}$, $\lambda_{5}$), as these factors are crucial for dexterous manipulation and contact-rich pinching tasks. The weight of 0.1 for the wrist orientation term ($\lambda_{2}$) is relatively smaller as the objective aims to emphasize fingertip tracking accuracy rather than wrist pose accuracy. The weight for the fingertip position term ($\lambda_{3}$) is set to 1 so that its contribution is of the same order of magnitude as the other terms when evaluated in typical retargeting scenarios.
For the weights of joint position and velocity regularization terms $w_{j}^{\text{joint}}$ and $w_{j}^{\text{vel}}$, index $j$ from 0 to 6 corresponds to the joints of the Panda arm, while indices $j=7$ to $22$ and $j=7$ to $30$ correspond to the joints of the Leap hand and the Shadow Hand respectively. Note that here we assume all DoFs of the Shadow Hand are actuated.
The joint position regularization weights $w_{j}^{\text{pos}}$ are determined according to the physical roles of different joints. For Leap hand, $j=7,11,15$ correspond to the abduction/adduction joints of index, middle and ring, $j=18$ corresponds to the DIP joint of the ring, and $j=20$ corresponds to the rotation of the thumb. For Shadow Hand, $j=9,13,17,22$ correspond to the finger movements of index, middle, ring and little finger in the palm plane, while $j=26$ corresponds to the rotation of the thumb. Non-zero weights are assigned to these joints whose extreme values tend to produce unnatural or mechanically unfavorable configurations. For joints with non-zero position regularization, the pre-defined joint configurations are set to $\bar{q}_{j}=0$.
The joint velocity regularization weights $w_{j}^{\text{vel}}$ are chosen to penalize large changes in joint positions compared to previous timestep to encourage trajectory smoothness. In practice, we use larger $w_{j}^{\text{vel}}$ for the arm joints and slightly smaller values for the finger joints, as arm motions are more prone to sudden changes. As shown in Fig. 7, this choice effectively reduces acceleration spikes and leads to smoother joint trajectories.
In our implementation, we rescale the size of the human hand by a factor of 1.5 for the Leap hand and 1.0 for the Shadow Hand to address the size difference between human and robot hand. In the real-world experiments, the retargeting control frequency is 20 Hz, and we use an exponential moving average with $\alpha_{\rm ema}=0.3$ to further smoothen the joint movements:
$\bm{q}_{t}=\alpha_{\rm ema}\cdot\bm{q}_{t}+(1-\alpha_{\rm ema})\cdot\bm{q}_{t-1}$ (7)
Discussion on Runtime, Latency, and Robustness
Runtime and latency
We profiled the computational cost of our retargeting pipeline on the LEAP-hand kinematic retargeting experiments. The full teleoperation loop runs at 20 Hz, corresponding to a frame time of 50 ms. Within each frame, the nonlinear optimizer for the retargeting objective takes on average 33 ms, which is lower than the frame time and is suitable for real-time control.
Overall, latency in the system mainly comes from three stages: (i) hand pose estimation and streaming from the Vision Pro headset; (ii) retargeting optimization; (iii) communication latency between our teleoperation node and the robot; and (iv) the low-level control latency on the robot side. The Vision Pro perception module runs in a separate thread and continuously receives hand pose estimations, while our 20 Hz teleoperation loop always uses the latest available hand pose at the beginning of each control cycle. After that, the optimizer runs within about 33 ms and the resulting joint commands are immediately sent to the low-level joint controllers. The latency introduced by our retargeting module is below 50 ms, and the Vision Pro streaming introduces an additional latency of about 50 ms[park2024avp]. Overall, the end-to-end latency of the system is approximately 0.15–0.25 s, which we empirically found sufficient for ensuring real-time teleoperation in all our real-world manipulation tasks.
Robustness to tracking noise and occlusions
Accurate detection and tracking of human hand pose is key to dexterous teleoperation. We experimented with both a single RGB camera and the Vision Pro headset for hand pose detection. Compared to the RGB camera, Vision Pro provides a more flexible field of view and more accurate hand pose tracking, but it still inevitably exhibits some tracking noise, especially in precise motions such as pinching. In practice, we observed that when the human thumb and index fingertip are already in contact, the estimated fingertip distance from Vision Pro can remain slightly positive and exhibit small jitter over time. We specially design the pinch term in our objective to be more robust to tracking error to enable stable pinching: we rescale the estimated fingertip distances in a small range $\epsilon_{2},\epsilon_{1}]$ to $0,\epsilon_{1}]$ and clamp values below $\epsilon_{2}$ to zero (Sec. III-A), so that sufficiently small estimated distances are treated as a closed pinch.
With respect to occlusions, Vision Pro can estimate a consistent hand pose under slight self-occlusion conditions, such as when a finger is partially occluded by the palm, as long as the operator keeps the hand within the headset’s field of view. This was sufficient for all three teleoperation tasks in our experiments, but for scenarios that require robustness under severe occlusions, Vision Pro may produce inaccurate hand pose estimation. Glove-based hand tracking systems will offer better robustness and be more occlusion-resistant in such cases[wang2024dexcap, zhang2025doglove].